OpenStudy (anonymous):

Let a, b, c, and n be positive integers. If a + b + c = 19 *97 and a+n=b-n=c/n; compute value of a

OpenStudy (rational):

a+n = c/n b-n = c/n ------------ a+b = 2c/n a+b+c = 19*97 2c/n + c = 19*97 (n+2)*c/n = 19*97

OpenStudy (rational):

Next notice that the right hand side is a product of two "primes", so there are only finite number of combinations for the left hand side product. Work them by cases

OpenStudy (rational):

letting n+2 = 19 and c/n = 97 gives you a = 80

OpenStudy (anonymous):

n=17 then

OpenStudy (rational):

coincidence the answer i got in below question is also 80 ! http://openstudy.com/study#/updates/55046957e4b079ad354fc2af